# Latency-Response Theory Model: Evaluating Large Language Models via Response Accuracy and Chain-of-Thought Length

### v1

@@ Abstract
The proliferation of Large Language Models (LLMs) necessitates valid evaluation methods to provide guidance for both downstream applications and actionable future improvements. The Item Response Theory (IRT) model with Computerized Adaptive Testing has recently emerged as a promising framework for evaluating LLMs via their response accuracy. Beyond simple response accuracy, LLMs’ chain of thought (CoT) lengths serve as a vital indicator of their reasoning ability. To leverage the CoT length information to assist the evaluation of LLMs, we propose the Latency-Response Theory (LaRT) model, which jointly models both the response accuracy and CoT length by introducing a key correlation parameter between the latent ability and the latent speed. We derive an efficient stochastic approximation Expectation-Maximization algorithm for parameter estimation. We establish rigorous identifiability results for the latent ability and latent speed parameters to ensure the statistical validity of their estimation. Through both theoretical asymptotic analyses and simulation studies, we demonstrate LaRT’s advantages over IRT in terms of superior estimation accuracy and shorter confidence intervals for latent trait estimation. To evaluate LaRT in real data, we collect responses from diverse LLMs on popular benchmark datasets. We find that LaRT yields different LLM rankings than IRT and outperforms IRT across multiple key evaluation metrics including predictive power, item efficiency, ranking validity, and LLM evaluation efficiency. Code and data are available at https://github.com/Toby-X/Latency-Response-Theory-Model.
@@ 1 Introduction
Third, to evaluate LaRT in practical settings, we conduct comprehensive empirical data analysis on real data by collecting responses from diverse LLMs on popular benchmark datasets. We observe an interpretable and consistent trend across all benchmark datasets: LLMs with greater latent ability exhibit lower latent speed (longer CoT). Quantitatively, we compare the performance of LaRT against IRT on the real-world data. We find that LaRT yields different model rankings than IRT. Moreover, LaRT outperforms IRT across multiple key evaluation metrics in LLM evaluation: predictive power (prediction accuracy on unseen question items), validity (ranking consistency over different benchmark datasets), item efficiency (the number of items required for evaluation), and LLM efficiency (the number of LLMs required for evaluation).
@@ 3 Estimation Algorithm
While SAEM has been widely used in latent variable modeling, its successful application to LaRT requires three additional model-specific components that tailor implementation strategies to ensure computational efficiency and algorithmic stability: (1) an efficient sampler for the latent variables that can substantially improve both computational speed and convergence behavior, (2) a suitable initialization strategy that can reduce the number of iterations and enhance estimation accuracy, and (3) a method or estimating individual LLM latent variables after obtaining population parameters. his section develops each component in turn. We first propose an efficient SAEM algorithm for estimating the population parameters of LaRT in Subsection 3.1, addressing the aspects discussed above. Next, in Subsection 3.2, we introduce an effective spectral-based initialization method for the proposed SAEM. Finally, in Subsection 3.3, we describe how to estimate the latent variables of LLMs once the population parameters have been obtained, under both static and active evaluation settings.
@@ 3.2 Initialization for the SAEM algorithm
Effective initialization is critical for SAEM’s efficiency and stability. We present a spectral-based initialization that is non-iterative, data-driven, and statistically consistent. This approach exploits the fact that the probit link for response accuracy and the log-normal specification for CoT fall within the scope of generalized linear factor models (GLMs) or linear factor models, allowing us to adapt the spectral-based method of zhang2020note, originally developed for GLM settings, to initialize the LaRT framework. Specifically, we extend their procedure—which was designed for single-modality response-accuracy data—to accommodate the bimodal structure of response accuracy and CoT, introducing additional steps to initialize the CoT-related parameters under the LaRT hierarchical model. Algorithm 2 summarizes the full algorithm.
@@
This informed initialization enables the algorithm to adopt a decaying step size $\alpha_{t}$ from the very first iteration. In contrast, standard SAEM implementations typically require an initial burn-in phase, during which a large step size (e.g., $\alpha_{t}=1$) must be used for many iterations (lavielle2014improved; kuhn2004coupling; camilli2019stochastic). This phase is crucial when using random initializations, which often start far from the optimum and therefore require large updates to move into the optimal region. However, it also introduces an additional tuning burden: the appropriate length of the burn-in phase varies with the initial values, making the procedure more unstable and sensitive to initialization. This initialization strategy bypasses this burn-in phase entirely by replacing random initialization with an efficient, non-iterative, data-driven strategy, while also enjoying favorable statistical consistency properties (zhang2020note). Our spectral initialization starts near the optimum, allowing SAEM to use a decaying step size from iteration one. This yields both faster convergence and improved estimation accuracy ompared with traditional SAEM implementations, as demonstrated in Appendix D.
3.3 Maximum-a-posterior estimation for individual latent ability and speed.
@@ 6 Applications to Real LLM Data
When collecting response data from LLMs, we set the maximum CoT length to 10,240 tokens. This limit is generous for mathematical reasoning, as ground truth solutions in the MATH dataset rarely exceed 1,000 tokens (hendrycksmath2021). By setting the limit to 10,240, we allow LLMs to utilize ten times the token budget of human references. Furthermore, this setting aligns with the “medium-to-high” reasoning regime defined by agarwal2025gpt, who evaluated gpt-oss across limits of 1,000, 6,000, and 16,000 tokens. Detailed hyperparameter configurations are in Appendix E.
After obtaining LLMs’ responses to each math problem, we use each LLM’s tokenizer to count the length of their CoT. We delete LLMs who failed to answer any question correctly in these four benchmarks. After data preprocessing, there are 138 LLMs for further evaluation.
Given this collected dataset of LLM responses, we apply LaRT to evaluate LLMs on these representative math benchmarks. First, we visualize the evaluation results based on the estimated latent ability and latent speed to assess consistency and demonstrate how LaRT automatically adjusts rankings by accounting for item characteristics across benchmarks. A key finding is strong negative correlation between ability and speed: higher-ability LLMs produce longer CoT. Second, we examine the benchmark characteristics through the estimated population parameters, highlighting their difficulty levels and discriminative power. We also compare LaRT and IRT across four key evaluation metrics from the LLM evaluation literature (Subsection 6.2) and quantitatively demonstrate LaRT’s superiority over IRT across all four (Subsection 6.3).

### v2

@@ Abstract
The proliferation of Large Language Models (LLMs) necessitates valid evaluation methods to guide downstream applications and actionable future improvements. The Item Response Theory (IRT) has recently emerged as a promising framework for evaluating LLMs via their response accuracy. Beyond simple response accuracy, LLMs’ chain of thought (CoT) lengths serve as a vital indicator of their reasoning ability. To leverage the CoT length information to assist the evaluation of LLMs, we propose Latency-Response Theory (LaRT) to jointly model the response accuracy and CoT length by introducing the latent ability, latent speed, and a key correlation parameter between them. We derive an efficient estimation algorithm and establish rigorous identifiability results for the population parameters to ensure the statistical validity of estimation. Theoretical asymptotic analyses and simulation studies demonstrate LaRT’s advantages over IRT in terms of higher estimation accuracy and shorter confidence intervals for latent traits. A key finding is that the asymptotic estimation precision of the latent ability under LaRT exceeds that of IRT whenever the latent ability and latent speed are correlated. We collect real responses from diverse LLMs on popular benchmark datasets. The application of LaRT reveals a strong negative correlation between the latent ability and latent speed in all benchmarks, with stronger correlation for more difficult benchmarks. This finding supports the intuition that higher reasoning ability correlates with slower speed and longer response latency. LaRT yields different LLM rankings than IRT and outperforms IRT across multiple key evaluation metrics including predictive power, item efficiency, ranking validity, and LLM evaluation efficiency. Code and data are available at https://github.com/Toby-X/Latency-Response-Theory-Model.
@@ 1 Introduction
Third, to evaluate LaRT in practical settings, we conduct comprehensive empirical data analysis on real responses collected from diverse LLMs on popular benchmark datasets: MATH500, AMC23, AIME24, and AIME25. We observe an interpretable and consistent trend across all benchmark datasets: LLMs with higher latent ability exhibit slower speed and longer response latency. This correlation is stronger for more difficult benchmark datasets. Quantitatively, we compare the performance of LaRT against IRT on the real-world data. We find that LaRT yields different model rankings than IRT. Moreover, LaRT outperforms IRT across multiple key evaluation metrics in LLM evaluation: predictive power (prediction accuracy on unseen question items), validity (ranking consistency over different benchmark datasets), item efficiency (the number of items required for evaluation), and LLM efficiency (the number of LLMs required for evaluation).
@@ 3 Estimation Algorithm
While SAEM has been widely used in latent variable modeling, its successful application to LaRT requires three additional model-specific components that tailor implementation strategies to ensure computational efficiency and algorithmic stability: (1) an efficient sampler for the latent variables that can substantially improve both computational speed and convergence behavior, (2) a suitable initialization strategy that can reduce the number of iterations and enhance estimation accuracy, and (3) a method or estimating individual LLM latent variables after obtaining population parameters. This section develops each component in turn. We first propose an efficient SAEM algorithm for estimating the population parameters of LaRT in Subsection 3.1, addressing the aspects discussed above. Next, in Subsection 3.2, we introduce an effective spectral-based initialization method for the proposed SAEM. Finally, in Subsection 3.3, we describe how to estimate the latent variables of LLMs once the population parameters have been obtained, under both static and active evaluation settings.
@@ 3.2 Initialization for the SAEM algorithm
Effective initialization is critical for SAEM’s efficiency and stability. We present a spectral-based initialization that is non-iterative, data-driven, and statistically consistent. This approach exploits the fact that the probit link for response accuracy and the log-normal specification for CoT fall within the scope of generalized linear factor models (GLFMs) or linear factor models, allowing us to adapt the spectral-based method of zhang2020note, originally developed for GLFM settings, to initialize the LaRT framework. Specifically, we extend their procedure—which was designed for single-modality response-accuracy data—to accommodate the bimodal structure of response accuracy and CoT, introducing additional steps to initialize the CoT-related parameters under the LaRT hierarchical model. Algorithm 2 summarizes the full algorithm.
@@
This informed initialization enables the algorithm to adopt a decaying step size $\alpha_{t}$ from the very first iteration. In contrast, standard SAEM implementations typically require an initial burn-in phase, during which a large step size (e.g., $\alpha_{t}=1$) must be used for many iterations (lavielle2014improved; kuhn2004coupling; camilli2019stochastic). This phase is crucial when using random initializations, which often start far from the optimum and therefore require large updates to move into the optimal region. However, it also introduces an additional tuning burden: the appropriate length of the burn-in phase varies with the initial values, making the procedure more unstable and sensitive to initialization. This initialization strategy bypasses this burn-in phase entirely by replacing random initialization with an efficient, non-iterative, data-driven strategy, while also enjoying favorable statistical consistency properties (zhang2020note). Our spectral initialization starts near the optimum, allowing SAEM to use a decaying step size from iteration one. This yields both faster convergence and improved estimation accuracy compared with traditional SAEM implementations, as demonstrated in Appendix D.
3.3 Maximum-a-posterior estimation for individual latent ability and speed
@@ 6 Applications to Real LLM Data
When collecting response data from LLMs, we set the maximum CoT length to 10,240 tokens. This limit is generous for mathematical reasoning, as ground truth solutions in the MATH dataset rarely exceed 1,000 tokens (hendrycksmath2021). By setting the limit to 10,240, we allow LLMs to utilize ten times the token budget of human references. Furthermore, this setting aligns with the “medium-to-high” reasoning regime defined by agarwal2025gpt, who evaluated gpt-oss across limits of 1,000, 6,000, and 16,000 tokens. Detailed hyperparameter configurations are in Appendix E. After obtaining LLMs’ responses to each math problem, we use each LLM’s tokenizer to count the length of their CoT. We delete LLMs who failed to answer any question correctly in these four benchmarks. After data preprocessing, there are 138 LLMs for further evaluation.